how long would a laser have to be on to emit only one photon? Say the power of the laser is 1 W at 500 nm.
Answer #1

Given , the power of laser is 1 W or J/s.

i.e. in 1 second laser is emitting photons of energy 1 Joule.

The wavelength (lambda) of laser = 500 nm = 500 x 10-9 m

We know that

  E = rac{hC}{lambda}

Where, h = 6.63 x 10-34 J-s is Planck's constant

E is energy of the photon

C = 3 x 108 m/s is speed of the light

So, energy of a single photon emitting from the laser is  

6.63 × 10-34 × 3 × 108 500 × 10-9 ^ ^ E= = 3.978 × 10-19 Joule

According to the question, the laser is emitting 1 Joule of energy per second.

So, total number of photons (N) emitting from the laser in 1 second is :

N = rac{1}{3.9 imes 10^{-19}} = 2.5 imes 10^{18}

I.e. time taken by laser to emit 2.5 x 1018 photons = 1 second

So, time taken by laser to emit one photon = 1 / (2.5 x 1018) = 3.978 x 10-19 second

Hence, the laser have to be on for 3.978 x 10-19 second to emit only one photon.

For any doubt please comment and please give an up vote. Thank you.